# XFormParser: A Simple and Effective Multimodal Multilingual Semi-structured Form Parser

### v1

@@ Abstract.
In the domain of document AI, semi-structured form parsing plays a crucial role and is employed across a broad spectrum of academic and industrial applications, including structured data management, knowledge retrieval, and content generation. This task encompasses a complex form of data mining that leverages techniques from key information extraction (KIE), dealing with inputs that range from plain text to intricate modal data comprising images and structural layouts. The advent of pre-trained multimodal models significantly empowers document AI frameworks to adeptly extract pivotal information from form documents presented in diverse formats such as PDFs and images. Nonetheless, the endeavor of form parsing is still encumbered by notable challenges like subpar capabilities in multi-lingual parsing and diminished recall in contexts rich in text and visuals. In this work, we introduce a simple but effective Multimodal and Multilingual semi-structured FORM PARSER (XFormParser), which anchored on a comprehensive pre-trained language model and innovatively amalgamates semantic entity recognition (SER) and relation extraction (RE) into a unified framework, enhanced by a novel staged warm-up training approach that employs soft labels to significantly refine form parsing accuracy without amplifying inference overhead. Furthermore, we have developed a groundbreaking benchmark dataset, named InDFormBench, catering specifically to the parsing requirements of multilingual forms in various industrial contexts. Through rigorous testing on established multilingual benchmarks and InDFormBench, XFormParser has demonstrated its unparalleled efficacy, notably surpassing the state-of-the-art (SOTA) models in RE tasks within language-specific setups by achieving an F1 score improvement of up to 1.79%. Our framework exhibits exceptionally improved performance across tasks in both multi-language and zero-shot contexts when compared to existing SOTA benchmarks. The robustness and reliability of XFormParser are thoroughly validated through comprehensive evaluations on the InDFormBench benchmark. [^1] The codes, datasets, and pre-trained models will be released.
Semi-structured form, Information extraction, Multi-modal, multilingual knowledge, Pre-trained language model
conference: ; ; submissionid: 286
1. Introduction
@@
Document AI is the technology of automatically reading, understanding, and analyzing business documents. It is widely applied in commercial, governmental, and educational sectors and is crucial to departmental efficiency and productivity. A key task of Document AI is to parse and extract form information from scanned documents. Form parsing is essentially an entity relation mining task that connects the Named Entity Recognition (NER) task (Li et al., 2020) and Key Information Extraction (KIE) task (Cui et al., 2021; Yu et al., 2021; Hong et al., 2022). Due to the diversity of the layout and the format, poor quality of scanned document images, and complexity of template structures, representing and understanding the unstructured information in documents using generic rules becomes a highly challenging task.
Figure 1. An illustration of named entity recognition for unstructured forms.
Unlike traditional NER tasks (Schuster et al., 2013; Esser et al., 2012) which only deal with textual information, and traditional pattern recognition tasks (Dengel and Klein, 2002; Medvet et al., 2011, 2011), form parsing requires consideration of multiple modalities. Apart from textual information, the position and layout of text blocks also play a crucial role in semantic interpretation. Previously, rules and workflows for each type of document often needed to be hard, requiring updates with changes in specific formats or when dealing with multiple formats. Current researchers believe that effective language models for form parsing need to comprehend target entities and adapt to different document formats in contexts involving multiple modalities. To address these issues, advanced Form Document AI models and algorithms are expected to automatically classify, extract, and structure information from business documents, minimizing manual intervention.
@@
Mainstream document intelligence methods typically involve using deep neural networks to model elements in documents from the perspectives of computer vision (He et al., 2016), natural language processing (Devlin et al., 2018), or Multimodal fusion (Atrey et al., 2010). Form parsing algorithms based on deep learning initially involved detecting and classifying specific regions of document images. Then, it utilized Optical Character Recognition (OCR) models (Li et al., 2023) to extract text information. This process is called unstructured to semi-structured, and the resulting form is called semi-structured. Subsequently, different categories of form blocks and their corresponding text content were routed to dedicated information extraction modules, constructing a pipeline to process entire-page documents(Xu et al., 2020b; Wang et al., 2022).
With the development of pre-training techniques (Zoph et al., 2020) and multimodal fusion technologies, researchers have developed a plethora of semi-structured form parsing algorithms based on Transformer (Vaswani et al., 2017). (Yu et al., 2021; Zhang et al., 2020) proposed a graph convolutional neural network information extraction model that fuses textual and visual information for form parsing in specific business scenarios. (Xu et al., 2020a) introduced the LayoutLM pre-trained model, which combined language models and Transformer, expanding BERT (Devlin et al., 2018) architecture by incorporating layout information to consider spatial relationships between tokens and textual content connections, resulting in outstanding performance in form parsing tasks. Furthermore, LayoutLMv2 (Xu et al., 2020b) and LayoutLMv3 (Huang et al., 2022) enhance the integration of multimodal information in the pretraining stage. They also add pretraining strategies for text-image alignment and text-image matching, incorporating word token alignment targets. LayoutLMv3 is the first multimodal model in Document AI that does not rely on pre-trained CNN (He et al., 2016) or Faster R-CNN (Girshick, 2015) backbones to extract visual features, further promoting the training efficiency and generalizability of the model. The XFormParser(Chen et al., 2023) model proposes a Global Structure Knowledge-guided relation extraction framework. The loop uses the ”generate-capture-merge” operation so that entity representation and global structural knowledge reinforce each other.
@@
Although these models have made significant progress in the field of Semi-structured form parsing, as research and applications have become more widespread, most existing methods still face three limitations: 1) The accuracy of key information extraction from complex, multilingual form images remains insufficient; 2) Multimodal models dominated by visual modalities still lag behind text-dominated multimodal models in form parsing tasks in rich-text and long-text scenarios; 3) General Multi-modal Large Language Models (MLLM), like GPT4 (Huang et al., 2023) and their training data are not completely open-source, making it impossible to directly access advanced pre-trained weights and tracks. Therefore, based on the self-supervised pre-training model, it is crucial to further mine the representation methods of entities and the relationships between entities in the background of multi-modal and multi-lingual knowledge and to study more effective fine-tuning paradigms in complex scenes for understanding the mechanism of table information extraction and effectively applying it to the industrial field.
@@
To address these issues, we propose a semi-structured form parser with multimodal and multilingual knowledge, named XFormParser. For input data from semi-structured forms, XFormParser utilizes a text-centric pre-trained model to generate vectors containing text information, visual information, and spatial positional information. Subsequently, these vectors are fed into the downstream joint network to complete two tasks: Semantic Entity Recognition (SER) and Relation Extraction (RE), namely NER and KIE tasks for semi-structured form, to realize form parsing. We use the multilingual document understanding pre-trained model LayoutXLM (Xu et al., 2022) as the backbone of XFormParser. The SER Task obtains text box classification through fully connected layers, and the RE task learns the categories of entity relations through a decoder based on Bi-LSTM (Chung et al., 2014; Sun et al., 2022) and Biaffine (Nguyen and Verspoor, 2019). In addition, we further build InDFormBench, a Chinese and English multi-scenario form parsing dataset for industrial applications, based on public benchmarks. Training the model on this dataset helps XFormParser learn semi-structured forms from the real world and achieve new SOTA performance.
@@
  • We propose XFormParser, an efficient model for multi-language semi-structured form parsing based on a pre-trained Transformer. It integrates two tasks: SER and RE, along with the joint loss function and training method of soft labels warm-up in stages. XFormParser effectively enhances form parsing performance without additional inference resources and overhead.
@@
  • We construct InDFormBench, a cross-scenario form parsing benchmark dataset in both Chinese and English. It contains 500 form images collected from eight major industrial application scenarios, along with corresponding annotation information in JSON format semi-automatically generated and rigorously validated using GPT4.
@@
  • Through experiments and analysis, we confirm that XFormParser achieves the F1 score of at most 1.79% over the SOTA model for RE tasks in Language-specific scenarios. XFormParser achieves significantly better results than SOTA for both dual-task in Multi-Language and Zero-shot scenarios. We also verify the effectiveness and reliability of XFormParser through experiments on InDFormBench.
2. Related Work
2.1. Single Modality-Based Extraction Methods.
In recent years, various document AI tasks based on natural language understanding (NLU) or computer vision (CV) have made great progress (alm; Yang et al., 2022c, a; Mo et al., 2024; Cheng et al., 2024; Yang et al., 2024; Wang et al., 2023; Ro et al., 2020). CloudScan (Palm et al., 2017) attempts to use LSTM to extract information from invoice-type form documents, on this basis, Sage et al. (Sage et al., 2019) added contextual semantic and spatial information for information extraction. Nguyen et al. (Nguyen and Verspoor, 2019), obtained the input vector by concatenation of word embedding and character embedding, entered it into the LSTM-CRF-based NER network to obtain the entity recognition results, and then obtained the relationship between entities through affine attention and softmax. REBEL (Cabot and Navigli, 2021) uses BART as a prototype to transform triplet extraction into a sequence-generating sequence problem through grammar rules. Chargrid (Katti et al., 2018) uses the OCR model of character granularity to extract characters, and CNN of the encoder-decoder is used for text box detection and semantic segmentation of key information for document images. Katti et al. (Katti et al., 2018) proposed using word2vec as a word-level feature extractor and fusing visual information to improve the extraction performance. Bertgrid (Denk and Reisswig, 2019), inspired by Chargrid (Katti et al., 2018), uses the encoder-decoder of a word-granularity convolutional neural network for semantic segmentation. ViBERTgrid (Lin et al., 2021) fuses text features from BERT with image features from convolutional neural networks to achieve better results.
2.2. Multilingual Multi-modal Fusion-based Methods.
The task of form information extraction has developed into a multi-modal task that can integrate CV and NLU (alm; Yang et al., 2024, 2020, 2022b, 2019; Chai et al., 2024), which needs to use the structure information and content information of the form. Most current multimodal tasks on form parsing are extensions of related CV and NLP tasks. The main challenge in form parsing is how to integrate visual language models, Examples such as Vision Transformer (ViT) (Wu et al., 2021) and contrast Language-Image Pretraining CLIP (Radford et al., 2019) have demonstrated the potential to bridge the gap between text and images. These models provide new ways to understand visual elements in the context of textual content.
@@ 2. Related Work
Figure 2. Overall architecture of the proposed XFormParser. For Multimodal input, XFormParser utilizes layoutXLM to generate vectors containing text, visual, and spatial positional information. Subsequently, these vectors are fed into the downstream joint network to complete SER and RE tasks. The SER Task obtains text box classification through fully connected layers, and the RE task learns the categories of entity relations through a decoder based on Bi-LSTM and Biaffine.
DocStruct (Wang et al., 2020) integrates visual, layout, and semantic information, utilizing a multilayer perceptron and optical character detection to extract layout and visual information. It uses visual features as mobile features and employs BERT to obtain semantic information. An attention mechanism gate is used to decide whether to use visual features as mobile information. Negative sampling is employed in the training method, achieving hierarchical extraction for form parsing. The effectiveness of key information extraction is validated on the FUNSD (Guillaume Jaume, 2019) dataset. TRIE (Zhang et al., 2020) combines text detection and information extraction, allowing the two tasks to mutually enhance each other for better information extraction results. The StructLM (Li et al., 2021) model learns cell layouts and understands which words belong to the same cell. Its pre-training tasks also learn the positional relationships at the text box level. The LiLT (Wang et al., 2022) model proposes a language-independent transformer focusing on text layout, introducing a new Bidirectional Attention Complement Mechanism (BiACM) to enhance cross-modal cooperation. It also proposes Key Point Localization (KPL) and Cross-modal Alignment Identification (CAI) tasks, combined with the widely used Masked Visual Language Model (MVLM) as pre-training objectives. ERNIE-Layout (Peng et al., 2022), proposed by Baidu, enhances layout knowledge throughout the workflow to learn better representations that combine text, layout, and image features. It reorders the input sequence during serialization and introduces related pre-training tasks, such as reading order prediction, to learn the correct reading order of documents. To improve the model’s layout awareness, spatial disentangled attention is integrated into the multimodal transformer, and the region replacement prediction task is incorporated into the pre-training phase. LayoutLMv3 (Huang et al., 2022) also pre-trains with word-patch alignment objectives, learning cross-modal alignment by predicting whether the corresponding image patches of text words are masked. Its simple unified architecture and training objectives make LayoutLMv3 a general pre-training model suitable for both text-centric and image-centric document AI tasks. XFormParser (Chen et al., 2023) first generates initial relation predictions for entity pairs extracted from document scan images. It then captures global structure knowledge from previous iterative predictions and incorporates this knowledge into entity representations. This ”generate-capture-incorporate” loop is repeated multiple times, allowing entity representations and global structure knowledge to reinforce each other.
3. Method
3.1. Task Definition
@@
The semi-structured form after OCR processing consists of a list of semantic entities, each of which consists of a set of words, the coordinates of the bounding box, and a picture, called the cell of the form and defined as:
(1) $b_{i}=[[w_{1i},...,w_{mi}],[x_{1i},x_{2i},y_{1i},y_{2i}],[image]]$
@@
Where $[w_{1i},\dots,w_{mi}]$ represents the word group, $x_{1i}/x_{2i}$ and $y_{1i}/y_{2i}$ are the upper-left and lower-right coordinates respectively, and image represents the visual information. The documents in the dataset we used are annotated with labels for each entity and relationships between entities. We represent each comment document as:
(2) $D=[[{b_{1},\ldots,b_{n}}],[{l_{1},\ldots,l_{n}}],[{(b_{1},h_{1}),\ldots,(b_{m},b_{h_{m}})}]]$
@@
where $l$ is the label of each entity and $L$ is a predefined set of entity labels. $(bi,bhi)$ refers to the relationship between entities $b_{i}$ and $b_{hi}$, as well as the link point from $b_{hi}$ to $b_{i}$. It is worth noting that the entity may have relationships with multiple entities or may not have a relationship with any other entity.
@@
Cell Classification Task (SER). The cell classification problem needs to classify all the cells and obtain a list of resulting labels $L=[{l_{1},\ldots,l_{n}}]$, where $l_{i}$ corresponds to the cell labels, which can be: Header entity (HEADER), QUESTION entity (QUESTION), ANSWER entity (ANSWER) and OTHER entity (OTHER). When the entity classification is completed, each token is converted into a BIO tag according to the entity category.
@@
Relationship Extraction Task (RE). Given the above definition and description of a form, for a form D, each cell bi needs to find its corresponding sequence of cells in the whole form $B_{b_{i}}=[{b_{j},\ldots,b_{k}}]$, Finally, the normalized cell relation $R_{b_{i}}=[\{(b_{i},b_{j}),\ldots,$
$(b_{i},b_{n})\}]$ is obtained.
3.2. Overall Architecture

### v2

@@ Abstract
In the domain of Document AI, parsing semi-structured image form is a crucial Key Information Extraction (KIE) task. The advent of pre-trained multimodal models significantly empowers Document AI frameworks to extract key information from form documents in different formats such as PDF, Word, and images. Nonetheless, form parsing is still encumbered by notable challenges like subpar capabilities in multilingual parsing and diminished recall in industrial contexts in rich text and rich visuals. In this work, we introduce a simple but effective Multimodal and Multilingual semi-structured FORM PARSER (XFormParser), which anchored on a comprehensive Transformer-based pre-trained language model and innovatively amalgamates semantic entity recognition (SER) and relation extraction (RE) into a unified framework. Combined with Bi-LSTM, the performance of multilingual parsing is significantly improved. Furthermore, we develop InDFormSFT, a pioneering supervised fine-tuning (SFT) industrial dataset that specifically addresses the parsing needs of forms in various industrial contexts. XFormParser has demonstrated its unparalleled effectiveness and robustness through rigorous testing on established benchmarks. Compared to existing state-of-the-art (SOTA) models, XFormParser notably achieves up to 1.79% F1 score improvement on RE tasks in language-specific settings. It also exhibits exceptional cross-task performance improvements in multilingual and zero-shot settings.1
XFormParser: A Simple and Effective Multimodal Multilingual Semi-structured Form Parser
Xianfu Cheng1†, Hang Zhang1†, Jian Yang2, Xiang Li2, Weixiao Zhou1, Fei Liu3, Kui Wu1, Xiangyuan Guan1, Tao Sun1, Xianjie Wu1, Tongliang Li4*, Zhoujun Li1,5*, 1CCSE, Beihang University, 2Beihang University, 3Beijing Language and Culture University, 4Beijing Information Science and Technology University, 5Shenzhen Intelligent Strong Technology Co.,Ltd. {buaacxf, zhbuaa0, jiaya, xlggg, wxzhou, gxy0615, buaast, wuxianjie, lizj}@buaa.edu.cn, 17861517116@163.com, wukui0099@gmail.com, tonyliangli@bistu.edu.cn
†Both authors contributed equally to this research.* Corresponding Author.1 The codes, datasets, and pre-trained models are publicly available at https://github.com/zhbuaa0/xformparser.
1 Introduction
@@
Document AI is the technology of automatically reading, understanding, and analyzing business documents. It is widely applied in commercial, governmental, and educational sectors and is crucial to departmental efficiency and productivity. A key task of Document AI is to parse and extract form information from scanned documents. As shown in Figure 1, form parsing is essentially an entity relation mining task that connects the Named Entity Recognition (NER) task Li et al. (2020) and KIE task Cui et al. (2021); Yu et al. (2021); Hong et al. (2022). Due to the diversity of the layout and the format, poor quality of scanned document images, and complexity of template structures, representing and understanding the unstructured information in documents using generic rules becomes a highly challenging task.
Figure 1: An illustration of named entity recognition for unstructured forms.
Unlike traditional NER tasks which only deal with textual information, and traditional pattern recognition tasks Medvet et al. (2011); Cheng et al. (2020), mainstream Document AI methods typically involve using deep neural networks to model elements in documents from the perspectives of computer vision (CV) He et al. (2016); Ren et al. (2016), natural language processing (NLP) Zhou et al. (2023); Zhang et al. (2024c), or multimodal fusion Li et al. (2024). Apart from textual information, the position and layout of text blocks also play a crucial role in semantic interpretation.
@@
Form parsing algorithms based on deep learning initially involved detecting and classifying specific regions of document images. Then, it utilized Optical Character Recognition (OCR) models Li et al. (2023); Cheng et al. (2024) to extract text information. This process is called unstructured to semi-structured, and the resulting form is called semi-structured. Subsequently, different categories of form blocks and their corresponding text content were routed to dedicated information extraction modules, constructing a pipeline to process entire-page documents Xu et al. (2020b); Wang et al. (2022). Current researchers believe that effective language models for form parsing must comprehend target entities and adapt to different document formats in contexts involving multiple modalities. Advanced Document AI models Wang et al. (2020); Zhang et al. (2020); Li et al. (2021); Peng et al. (2022) are expected to automatically classify, extract, and structure information from business documents, minimizing manual intervention.
@@
Although advanced models have made significant progress in the field of Semi-structured form parsing, as research and applications have become more widespread, most existing methods still face three limitations: 1) The accuracy of key information extraction from complex, multilingual form images remains insufficient; 2) Multimodal models dominated by visual modalities still lag behind text-dominated multimodal models in form parsing tasks in rich-text and long-text scenarios; 3) Multi-modal large language models (MLLM) such as GPT4o Huang et al. (2023); Islam and Moushi (2024) and LayoutLLM Fujitake (2024) are difficult to deploy to the end side and achieve fast and high-performance inference via CPUs or low-memory GPUs due to the excessive weight of model parameters. Therefore, in industrial applications, it is a crucial research direction to further mine the entity classification and the relations between entities in multimodal and multilingual forms based on simple and effective pre-trained models (PTM) Zoph et al. (2020), and to study more effective fine-tuning paradigms in complex scenes.
@@
To address these issues, we propose a simple but effective semi-structured form parser with multimodal and multilingual knowledge, named XFormParser. For input data from semi-structured forms, XFormParser utilizes the multilingual document understanding PTM LayoutXLM Xu et al. (2022) to generate vectors containing text, visual, and spatial positional information. Subsequently, these vectors are fed into the downstream joint network to complete two tasks: Semantic Entity Recognition (SER) and Relation Extraction (RE), to realize form parsing. The SER Task obtains text box classification through fully connected layers, and the RE task learns the categories of entity relations through a decoder based on Bi-LSTM Sun et al. (2022) and Biaffine Nguyen and Verspoor (2019). In addition, we further build InDFormSFT, a Chinese and English multi-scenario form parsing SFT dataset for industrial applications, based on public benchmarks. Training the model on this dataset helps XFormParser learn semi-structured forms from the real world and achieve new SOTA performance.
@@
  • We propose XFormParser, which integrates two tasks: SER and RE, along with the joint loss function and training method of soft labels warm-up in stages. XFormParser effectively enhances form parsing performance without additional inference resources and overhead.
@@
  • We construct InDFormSFT, a cross-scenario form parsing SFT dataset in both Chinese and English. It contains 562 form images collected from 8 major industrial application scenarios and corresponding annotation information in JSON format that is semi-automatically generated using tools such as GPT4o and rigorously verified by humans.
@@
  • Through experiments and analysis, we confirm that XFormParser achieves an F1 score of at most 1.79% over the SOTA model for RE tasks in Language-specific scenarios. XFormParser achieves significantly better results than SOTA for both dual-task in Multi-language and Zero-shot scenarios. The ablation experiments on InDFormSFT demonstrate the effectiveness and robustness of XFormParser.
@@
Figure 2: Overall architecture of the proposed XFormParser. For Multimodal input, XFormParser utilizes layoutXLM to generate vectors containing text, visual, and spatial positional information. Subsequently, these vectors are fed into the downstream joint network to complete SER and RE tasks. The SER Task obtains text box classification through fully connected layers, and the RE task learns the categories of entity relations through a decoder based on Bi-LSTM and Biaffine.
2 Related Work
Limited by the lack of training data and the complexity of the corpus, relatively few effective models for parsing multilingual forms have been proposed in the past few years. Most applications use a pipeline approach to process form information one modality by one, such as XLM-RoBERTa Conneau et al. (2020), which is a multilingual version of RoBERTa Liu et al. (2019), and InfoXLM Chi et al. (2021), a multilingual pre-trained model Vaswani et al. (2017); OpenAI (2023) that maximizes the mutual information between multilingual and multi-granularity texts, which are used as text information processors to analyze forms.
The Visual-Language models Wu et al. (2021); Radford et al. (2019); Zhang et al. (2024b) demonstrate the potential to effectively align gaps between textual and other modal features, which can be used to integrate form structure information and content information. It is further revealed that form information extraction can be achieved by multi-modal techniques combining CV Cheng et al. (2022) and NLP Yang et al. (2024, 2020, 2022, 2019); Chai et al. (2024); Zhang et al. (2024a).
The LiLT Wang et al. (2022) model proposes a language-independent transformer focusing on text layout, introducing a new Bidirectional Attention Complement Mechanism to enhance cross-modal cooperation. It also proposes Key Point Localization and Cross-modal Alignment Identification tasks, combined with the widely used Masked Visual Language Model as pre-training objectives. LayoutLM Xu et al. (2020a) introduced the PTM, which combined language models and Transformer, expanding BERT Devlin et al. (2018) architecture by incorporating layout information to consider spatial relations between tokens and textual content connections, resulting in outstanding performance in form parsing tasks. Furthermore, LayoutLMv2 Xu et al. (2020b) and LayoutLMv3 Huang et al. (2022) enhance the integration of multimodal information in the pretraining stage. They also add pretraining strategies for text-image alignment and text-image matching, incorporating word token alignment targets.
Regarding the joint modeling of SER and RE, existing work primarily targets sequence labeling tasks, training jointly at the token granularity Jiang et al. (2024); Ji et al. (2024); Wang et al. (2023); El Khbir et al. (2024); Liu et al. (2023). At present, there is a lack of available models or methods for joint pattern training with cell information.
GeoLayoutLM Luo et al. (2023) improves the feature representation of text and layout by explicitly modeling geometric relationships and special pre-training tasks, which can improve the performance of form information extraction. However, the complex model structure makes the pre-training cost of related tasks high and too dependent on data. GOSE Chen et al. (2023) first generates initial relation predictions for entity pairs extracted from document scan images. It then captures global structure knowledge from previous iterative predictions and incorporates it into entity representations. This "generate-capture-incorporate" loop is repeated multiple times, allowing entity representations and global structure knowledge to reinforce each other.
3 Method
3.1 Task Definition
@@
The semi-structured form processed by the OCR model can be represented as a list of $n$ semantic entities. Each entity consists of a set of words named $Words_{i}$, the coordinates of the bounding box named $Bbox_{i}$, and an image named $Image_{i}$, called $i$-th cell of the form and defined as $b_{i}$:
$b_{i}=[Words_{i},Bbox_{i},Image_{i}]$ (1)
@@
The documents in our dataset are annotated with labels for each entity and relations between entities. We represent each comment document $D$ as:
$D=[B,L,R]$ (2)
@@
where $B$=$[b_{1},\ldots,b_{n}]$ denote all the cells of the form; $L$=$[{l_{1},\ldots,l_{n}}]$ is a predefined set of entity labels and $l$ is the label of each entity; $R$=$[{(b_{1},b_{2}),\ldots,(b_{j},b_{k})}]$ is the set of relations between entities, $(b_{j},b_{k})$ refers to the relation between $j$-th entity and $k$-th entity, as well as the link point from $b_{k}$ to $b_{j}$. It is worth noting that one entity may have relations with multiple entities or may not have a relationship with some other entity.
@@
Semantic Entity Recognition Task. The semantic entity recognition needs to classify all the cells and obtain a list of resulting labels $L$, the label of cells can be: Header entity (HEADER), QUESTION entity (QUESTION), ANSWER entity (ANSWER) and OTHER entity (OTHER). When the entity classification is completed, each token needs to be converted into a BIO label according to the entity category, to facilitate the alignment with experiments on open-source benchmark datasets.
@@
Relation Extraction Task. Given the above definition and description of a form, for a form $D$, each cell $b_{i}$ needs to find its corresponding sequence of cells in the whole form $B_{b_{i}}$=$[{b_{j},\ldots,b_{k}}]$, Finally, the normalized cell relation $R_{b_{i}}$=[$[\{(b_{i},b_{j}),\ldots,(b_{i},b_{k})\}]$ is obtained.
3.2 Overall Architecture